Church Inc. is presently enjoying relatively high growth because of a surge in the demand for its new product. Management expects earnings and dividends to grow at a rate of 37% for the next 4 years, after which competition will probably reduce the growth rate in earnings and dividends to zero, i.e., g = 0. The company's last dividend, D, was $1.25, its beta is 1.20, the market risk premium is 5.50%, and the risk-free rate is 3.00%. What is the current price of the common stock?
a. $39.57
b. $49.37
c. $40.80
d. $42.02
e. $43.25
1
Expert's answer
2012-11-29T08:15:27-0500
Today's price equal:
P = D*1.37^4*(1-g)/(k-g) = D*(1-g)/(rfr+beta(rm - rfr)-g)=
$4.4*1/(0.03+1.2(0.055))= $4.4/0.096 =$43.25

Answer: e) $43.25